A vessel in the shape of a spherical shell 758 mm in diameter, 18 mm shell thickness is completely filled with a
Fluid at atmospheric pressure. Additional fluid is then pumped in till the pressure increases by 7 N/mm. Find the volume of this additional fluid given that 1/m=0.25 and
E=2.1X 10
1
Expert's answer
2021-07-20T21:27:01-0400

"d = 758 mm"


"t = 18 mm"


"p = 7 \frac{N}{mm}"


"E = 2.1 \times 10^ 5 \frac{N}{mm}"


"\mu = 0.25"


Change in diameter "(\delta d)" :


"\delta d = \frac{pd^2}{4tE}[1-\frac{1}{m}]"


"\delta d = \frac{7\times 758^2}{4\times 18\times 2.1\times 10^5}[1-0.25]=0.1995 mm"


Change in volume (δv)


"e_v = 3\times \frac{\delta d}{d}"


"V = \frac{\pi}{6}\times d^3 = \frac{\pi}{6}\times 758^2 =300841mm^3"


"\frac{\delta v}{v} = 3\times \frac{\delta d}{d}"


"\delta d =3 \times 300814 \times \frac{0.1995}{758} =237.51 mm^3"